Solve the equation.
4(x + 6) = (4x + 24)

Answer by tutorcecilia(2152)

4(x + 6) = (4x + 24) [multiply using the distributive property]
4x+24=4x+24
4x-4x+24=4x-4x+24
24=24
There are an infinate number of solutions.